Please help.
Tim kicks a ball off the ground. After t seconds, its height, h (in feet), is given by the formula h = -16t^2 + 64t. What is the maximum height reached by the ball?
Thanks

Answer by stanbon(75887)

Tim kicks a ball off the ground. After t seconds, its height, h (in feet), is given by the formula h = -16t^2 + 64t. What is the maximum height reached by the ball?
------------------
You have a quadratic with a = -16 and b = 64
The max occurs when t = -b/2a = -64/(2*-16) = 2 second
---------------------
The height at t = 2 sec is as follows:
h(2) = -16(2)^2 + 64(2) = 64 ft
===========================
Cheers,
Stan H.